OpenStudy (anonymous):

f(x)=ab^x that go through (0,20) and (4,25)

OpenStudy (anonymous):

\[20 = ab^0 = a*1 = a\] so a is 20 and then for b you can go \[25 = 20*b^4\] which would become \[5/4 = b^4 \] so \[b = \sqrt[4]{5/4}\] so \[f(x) = 20 * \sqrt[4]{5/4}^x\]